OpenStudy (anonymous):

Can anyone please explain how to solve this? Given 1 hour = 60 minutes, 1 day = 24 hours, and 1 lunar month = 28 days, calculate the number of minutes in 1 lunar month. A. 51.4 minutes B. 403.20 minutes C. 40,320 minutes D. 11.2 minutes E. 70 minutes

OpenStudy (chmvijay):

1 hr = 60 minutes 24 hr =60 * 24 = Y minutes (find Y) but 24 hr =1 day 1 day has = y minutes 28 days have =Y * 28 = Z minutes for 28 days but 28 days = i lunar month =Z minutes find out Y and Z

OpenStudy (anonymous):

Thank you. So I calculate 60 * 24 = 1.440 and then I calculated 1.440 * 28 and I got 40,320. is this correct?

OpenStudy (anonymous):

@chmvijay

OpenStudy (rane):

the answer is C

OpenStudy (rane):

U R RIGHT @RH

OpenStudy (anonymous):

Thank you!!!

OpenStudy (rane):

any time :)

OpenStudy (chmvijay):

ur right :)sorry i was away from PC :P